# A Novel Hybrid Time-Varying Graph Neural Network For Traffic Flow Forecasting

### v3

@@ Abstract
Real-time and precise traffic flow prediction is vital for the efficiency of intelligent transportation systems. Traditional methods often employ graph neural networks (GNNs) with pre-defined graphs to describe spatial correlations among traffic nodes in urban road networks. However, these pre-defined graphs are limited by existing knowledge and graph generation methodologies, offering an incomplete picture of spatial correlations. While time-varying graphs based on data-driven learning have attempted to address these limitations, they still struggle with adequately capturing the inherent spatial correlations in traffic data. Moreover, most current methods for capturing dynamic temporal correlations rely on a unified calculation scheme using a temporal multi-head self-attention mechanism, which often leads to inaccuracies. In order to overcome these challenges, we have proposed a novel hybrid time-varying graph neural network (HTVGNN) for traffic flow prediction. Firstly, a novel time-aware multi-attention mechanism based on time-varying mask enhancement was reported to more accurately model the dynamic temporal dependencies among distinct traffic nodes in the traffic network. Secondly, we have proposed a novel graph learning strategy to concurrently learn both static and dynamic spatial associations between different traffic nodes in road networks. Meanwhile, in order to enhance the learning ability of time-varying graphs, a coupled graph learning mechanism was designed to couple the graphs learned at each time step. Finally, the effectiveness of the proposed method HTVGNN was demonstrated with four real data sets. Simulation results revealed that HTVGNN achieves superior prediction accuracy compared to the state of the art space-time graph neural network models. Additionally, the ablation experiment verifies that the coupled graph learning mechanism can effectively improve the long-term prediction performance of HTVGNN.
@@
unmanned vehicle, model predictive control, trajectory tracking.
@@ I Introduction
  • It is challenging to capture the spatial correlations between road network nodes without prior knowledge. Although some works try to create spatial structure maps based on existing knowledge [19, 20], these maps often inaccurately represent spatial correlations due to the limitations of their underlying assumptions and methods. Some other works [21, 22] have made strides in adaptive spatial modeling without fixed graphs, but they still predominantly use a global adaptive graph across all time steps, restricting their learning potential.
@@
To tackle the identified challenges, this paper introduces a novel time-aware multi-head attention mechanism enhanced by time-varying masks, aiming to boost the computational accuracy and temporal sensitivity of attention models. We present an efficient coupled time-varying graph convolutional recurrent network, designed to capture the spatio-temporal dynamics of traffic flows. This network leverages time-varying graphs for each time step, enhancing its ability to learn spatial dependencies and long-term trends in road network traffic. By integrating time-varying graphs across different time steps, our model effectively incorporates both short-term and long-term traffic patterns. Furthermore, we propose a dynamic graph convolution approach, which utilizes a combined topology and similarity matrix as a mask, enabling more precise modeling of the dynamic spatial correlations in traffic networks.
@@
  • A novel temporal perception multi-head self-attention mechanism enhanced by a time-varying mask, diverging from traditional approaches is proposed. This mechanism dynamically adjusts attention calculations based on the temporal characteristics of input data, enabling more accurate capture of temporal dependencies among traffic nodes. The time-varying mask consists of three components: a static mask embedding for correcting the multi-head self-attention mechanism, and a dynamic mask embedding to augment the time-awareness of the mechanism.
  • A new graph learning strategy is introduced to simultaneously address static and dynamic spatial correlations in traffic networks. To enhance learning of static spatial correlations, we employ both global and local time-varying embeddings, facilitating the understanding of spatial relationships at various time steps. Additionally, a coupling graph learning mechanism is implemented to integrate graphs learned at different time steps. For more precise capture of dynamic spatial correlations, we define a mask matrix, derived from topology and semantic matrices, to refine the dynamic graphs produced by our strategy.
  • Simulation on four real-world datasets is made to evaluate our proposed HTVGNN method. The results demonstrate the effectiveness of HTVGNN, showing higher prediction accuracy than the latest space-time graph neural network models. What’s more, an ablation study confirms that the integrated coupled graph learning mechanism significantly enhances the model’s long-term prediction capabilities.
@@ IV Methodology
The present study proposes a time trend-aware multi-head self-attention transformer and a graph convolution GRU-based traffic flow prediction model. The former effectively captures local temporal features, while the latter simultaneously extracts spatio-temporal features. Firstly, we introduce the fundamental structure of HTVGNN and illustrate it in Fig. 1. To better represent intricate correlations among nodes, we propose static and dynamic time-varying graphs that learn changing spatial correlations through trainable node embeddings. Subsequently, graph convolution GRU is employed to extract spatio-temporal features for accurate traffic flow prediction. Detailed model structures and mechanisms are presented in subsequent subsections.
@@ IV-A An enhanced temporal perception multi-head self-attention
where $h$ is the total amount of attention heads. $W$ is the final output projection matrix. However, in traditional multi-head attention mechanism, the Key vectors ($K_{i}$) will cause unnecessary interference [17]. In addition, since the learned matrix $W_{i}^{q}$, $W_{i}^{k}$, and $W_{i}^{v}$ remained unchanged after training, the score calculated by the traditional multi-head attention equation defined in Eq. (2) can not fully perceive the temporal trend of self-attention. Therefore, a time-varying mask embedding generator is designed. The key process of the generator finds the daily submask embedding $\tilde{P}^{D}$ and the weekly submask embedding $\tilde{P}^{W}$ of the time corresponding to the current traffic flow $X$. Then, based on an independent trainable embedding $\tilde{E}\in R^{N\times d_{m}}$, we have proposed an enhanced temporal perception multi-head self-attention. For notational simplicity, our proposed enhanced temporal perception multi-head self-attention (ETPMSA), as shown in Fig.2, can be calculated as follows in Eq.(4):
@@
where $\widehat{W}$ is a learnable parameter vector, $head_{i}^{ETP}$ denotes the $i^{th}$ self-attention head which can be calculted by the following Eq.(5) to Eq.(9).
@@
where $Q_{i}\in\ R^{T\times N\times d_{q}\ },K_{i}\in\ R^{T\times N\times d_{k}\ }$ and $V_{i}\in\ R^{T\times N\times d_{v}\ }$ are the Query, Key and Value of the traditional time multi-head attention as defined in equation (2), $M_{t}\in R^{N\times N}$ is mask matrix which is defined according to the size of the input data at the time step $t$. It should be noted that, $m_{t}\in R^{N\times d_{m}}$ is the mask embedding which composed of three submask embedding $\tilde{E}\in R^{N\times d_{m}}$ and $\tilde{P}^{D}_{t}\in R^{N\times d_{m}}$ , $\tilde{P}^{W}_{t}\in R^{N\times d_{m}}$. Submask embedding $\tilde{P}^{D}_{t}$ and $\tilde{P}^{W}_{t}$ are the daily submask embedding $\tilde{P}^{D}$ and weekly submask embedding $\tilde{P}^{W}$ at $t$. The submask embedding $\tilde{P}^{D}$ and $\tilde{P}^{W}$ are employed to enhance the temporal perception by updating the key $(K)$ of multi-heads self-attention mechanism.
@@ IV-B Position Encodings
Among them, $pos$ represents the position and $d$ represents the feature dimension, $D$ denotes the dimension of the model.
@@ IV-C1 Graph learning
where $E$ is a trainable node embedding, $\ e_{t}$ is a trainable bias of the node embedding $E$. Herein, in order to ensure the diversity of $A_{t}^{s}$ at different time step, we set different bias $e_{t}$ for different time step $t^{th}$ during the training process, $Relu()$ is used as the activation function.
@@
where $W_{t-T+2},....W_{t}$ are learnable parameter matrices.after completing the graph coupling process,The coupling graph $\tilde{A}_{t}^{s}$ is utilized as the ultimate static adjacency matrix for spatial correlation modeling at each time step.
@@
the dynamic adjacency matrix $\tilde{A}^{v}_{k_{(i,j)}}$ at the $k^{th}$ time step between node $i$ and node $j$ is defined as Eq.(13) to Eq.(16)
@@
Then,we think dynamic changes of spatial correlations occuring in the current time should belong to the actual existence and most active nodes. such as nodes which located in adjency matrix and DTW matrix.thus,we combine the adjency matrix and DTW matrix as the mask matrix of dynamic spatial correlations.Finally the dynamic adjacency matrix $\tilde{A}^{v}_{t}$ can be define as Eq.(16):
@@ IV-C2 Coupling time-varying graph convolution gated recurrent unit
where $H_{t}$ and $h_{t}$ are the input hidden embedding and output hidden embedding at the time step $t$, respectively. ||||| | represents the concatenation operation, ${z}_{t}$ and ${r}_{t}$ represent reset gate and update gate the time step $t$, respectively. $\mathcal{G}_{s}$ and $\mathcal{G}_{d}$ represent the static graph operation and dynamic graph operation respectively.$\theta_{z_{1}}$,$\theta_{z_{2}}$ and $\theta_{r_{1}}$,$\theta_{r_{2}}$ and  $\theta_{c_{1}}$,$\theta_{c_{2}}$ are learnable parameters in CTVGCRM.
@@
Dataset PEMSD3 PEMSD4 PEMSD7 PEMSD8
@@ V-A Datasets
We validate the predictive performance of HTVGNN by conducting experiments on four real-world traffic datasets, namely PeMSD3, PeMSD4, PeMSD7, and PeMSD8. These datasets were collected by the California Department of Motor Transportation Performance Measurement System (PeMS). Please refer to Table I for detailed information.
@@ V-B Evaluation Metrics
where $\hat{y}_{i}$ and $y_{i}$ represent the predicted value and real value of traffic flow at node $n$ respectively.
@@ V-D2 Model Settings
The experiments are conducted on a machine equipped with dual NVIDIA RTX3090 and 32GB of memory. HTVGNN is trained using Ubuntu 18.04, PyTorch 2.0.0, and the Adam optimizer [46]. The learning rate is set to 0.001, while the remaining hyperparameters of the model are presented in the table IV.
@@ V-E Performance Comparison
Tables II and III present the comparison results of 12 prediction models on four datasets. Overall, our HTVGNN achieved exceptional prediction accuracy across all 12 prediction strides. The deep learning-based approaches outperformed traditional statistical-based approaches (VAR), indicating their efficacy in modeling highly nonlinear traffic flows. As depicted in Tables II and III, these approaches exhibited suboptimal prediction performance. Moreover, methods that capture spatiotemporal dependencies generally surpassed those solely modeling temporal dependencies (LSTM), underscoring the significance of spatial dependencies. DCRNN and STGCN employ predefined graphs to model spatial dependencies in road networks; however, their performance heavily relies on the quality of these predefined graphs. AGCRN utilizes adaptive graphs to capture spatial dependencies, which typically enhances predictions compared to predefined graphs. Nevertheless, both predefined and adaptive graphs remain static during the prediction stage, disregarding dynamic changes in the road network. In contrast, models that dynamically model spatial dependencies (GMAN, PDFormer) significantly outperform other approaches based on static graph structures (DCRNN, STGCN, GWN). This can be attributed to their ability to learn dynamically changing graph structures and provide a broader representation space for modeling more complex spatiotemporal dependencies.Compared to other models, HTVGNN incorporates a time-varying mask matrix to enhance the multi-head attention mechanism for accurately capturing dynamic temporal correlations. Moreover, HTVGNN employs a dedicated coupled graph learning method tailored for multi-head attention, effectively improving its performance in long-term prediction tasks. Additionally, to better model dynamic spatial correlations in road networks, we define a mask based on topological and semantic matrices to optimize the dynamic time-varying graphs. Overall, our model outperforms existing state-of-the-art approaches on four datasets, achieving superior prediction accuracy.
@@ V-F Abliation study
$\mathbf{w/o\ TV}$: Replace the time-varying graph in CTVGCRM with a topological graph, emphasizing its pivotal role in capturing temporal dynamic correlations.
@@
The experiments were conducted multiple times, and the optimal MAE, RMSE, and MAPE values for each variant on PEMSD4 and PEMSD8 are presented in Table V.
TABLE V: Ablation experiments on PEMSD4 and PEMSD8.
@@ V-F2 Effect analysis of different types of graph structure in CTVGCRM and HTVGNN
$\mathbf{CTVGCRM}$: This is a coupling time-varying graph convolution gated recurrent network.
$\mathbf{CTVGCRM-sl}$:Let different time slices be learned with different embeddings, but the spatial correlations learned in different time slices are not coupled.
$\mathbf{CTVGCRM-ag}$: All time slices are learned using the same embedding, in which case the time-varying graph is the same as AGCRN.
$\mathbf{CTVGCRM-s}$:Replace coupled time-varying graphs with topological graphs.
@@
Subsequently, we assess the impact of our proposed RTP multi-head self-attention and coupling time graph on the overall model performance by leveraging the aforementioned optimal mask embedding and dropout techniques. We devise a series of five variants to conduct ablation experiments on our model using PEMS04 and PEMS08 datasets. These variants aim to highlight the significance of time-varying mask embedding and time-varying embedding in our approach, as described below:
@@
$\mathbf{w/o-tm}$: The time-varying mask embedding is excluded from the RTP multi-head attention, while CTVGCRM remains unchanged as HTVGNN.
$\mathbf{w/o-cg}$: Remove the coupling part of the time-varying graph.
$\mathbf{w/o-bc}$: The time-varying mask embedding is excluded from the RTP multi-head attention, and the coupling part of the time-varying graph is removed.
The results presented in Table VIII demonstrate the superiority of HTVGNN over both w/o-tm, highlighting the positive impact of incorporating time-varying mask embedding on enhancing self-attention accuracy. Furthermore, the superior performance of HTVGNN compared to w/o-cg and the superior performance of w/o-cg compared to w/o-bc indicates its advantage in capturing spatial correlations in transportation networks by leveraging coupling time-varying graphs.
@@
The results presented in Table IX unequivocally demonstrate the superiority of HTVGNN with coupling time-varying graphs over its counterpart without such coupling. This advantage becomes increasingly pronounced as the time step increases. Furthermore, Table IX reveals that HTVGNN outperforms HTVGNN-wc by 3.92%, 4.04%, and 4.32% in terms of MAE, RMSE, and MAPE on PEMS04 dataset at a temporal resolution of 60 minutes, respectively. Similarly, on PEMS08 dataset at the same temporal resolution, HTVGNN achieves improvements of 5.43%, 3.20%, and 4.57% for MAE, RMSE, and MAPE.

### v4

@@ Abstract
Real-time and precise traffic flow prediction is vital for the efficiency of intelligent transportation systems. Traditional methods often employ graph neural networks (GNNs) with pre-defined graphs to describe spatial correlations among traffic nodes in urban road networks. However, these pre-defined graphs are limited by existing knowledge and graph generation methodologies, offering an incomplete picture of spatial correlations. While time-varying graphs based on data-driven learning have attempted to address these limitations, they still struggle with adequately capturing the inherent spatial correlations in traffic data. Moreover, most current methods for capturing dynamic temporal correlations rely on a unified calculation scheme using a temporal multi-head self-attention mechanism, which at some level might leads to inaccuracies. In order to overcome these challenges, we have proposed a novel hybrid time-varying graph neural network (HTVGNN) for traffic flow prediction. Firstly, a novel enhanced temporal perception multi-head self-attention mechanism based on time-varying mask enhancement was reported to more accurately model the dynamic temporal dependencies among distinct traffic nodes in the traffic network. Secondly, we have proposed a novel graph learning strategy to concurrently learn both static and dynamic spatial associations between different traffic nodes in road networks. Meanwhile, in order to enhance the learning ability of time-varying graphs, a coupled graph learning mechanism was designed to couple the graphs learned at each time step. Finally, the effectiveness of the proposed method HTVGNN was demonstrated with four real data sets. Simulation results revealed that HTVGNN achieves superior prediction accuracy compared to the state of the art spatio-temporal graph neural network models. Additionally, the ablation experiment verifies that the coupled graph learning mechanism can effectively improve the long-term prediction performance of HTVGNN.
@@
Traffic flow forecasting, Graph neural network, Transformer, Multi-head self-attention.
@@ I Introduction
  • It is challenging to capture the spatial correlations between road network nodes without prior knowledge. Although some works try to create spatial structure graphs based on existing knowledge [19, 20], these graphs often inaccurately represent spatial correlations due to the limitations of their underlying assumptions and methods. Some other works [21, 22] have made strides in adaptive spatial modeling without fixed graphs, but they still predominantly use a global adaptive graph across all time steps, restricting their learning potential.
@@
To tackle the identified challenges, this paper proposed a novel hybrid time-varying graph neural network for traffic flow prediction. A novel enhanced temporal perception multi-head self-attention mechanism enhanced by time-varying masks was presented to boost the computational accuracy and temporal sensitivity of attention models. An efficient coupled time-varying graph convolutional recurrent network was designed to capture the spatio-temporal dynamics of traffic flows. This network leverages time-varying graphs for each time step, enhancing its ability to learn spatial dependencies and long-term trends in road network traffic. By integrating time-varying graphs across different time steps, our model effectively incorporates both short-term and long-term traffic patterns. Furthermore, we proposed a dynamic graph convolution approach, which utilized a combined topology and semantic matrix as a mask, enabling more precise modeling of the dynamic spatial correlations in traffic networks.
@@
  • A novel temporal perception multi-head self-attention mechanism enhanced by a time-varying mask, diverging from traditional approaches is proposed. This mechanism dynamically adjusts attention calculations based on the temporal characteristics of input data, enabling more accurate capture of temporal dependencies among traffic nodes. The time-varying mask consists of three components: a static mask embedding for correcting the multi-head self-attention mechanism, and two dynamic mask embeddings to augment the time-awareness of the mechanism.
  • A new graph learning strategy is reported to simultaneously address static and dynamic spatial correlations in traffic networks. To enhance learning of static spatial correlations, we employ both global and local time-varying embeddings, facilitating the understanding of spatial relationships at various time steps. Additionally, a coupling graph learning mechanism is implemented to integrate graphs learned at different time steps. For more precise capture of dynamic spatial correlations, we define a mask matrix, derived from topology and semantic matrices, to refine the dynamic graphs produced by our strategy.
  • Simulation on four real-world datasets is made to evaluate our proposed HTVGNN method. The results demonstrate the effectiveness of HTVGNN, showing higher prediction accuracy than the latest space-time graph neural network models. Furthermore, an ablation study confirms that the integrated coupled graph learning mechanism significantly enhances the model’s long-term prediction capabilities.
@@ IV Methodology
In order to more accurately capture the dynamic spatio-temporal correlations between different traffic nodes, we have proposed a novel hybrid time-varying graph neural network (HTVGNN) for traffic flow prediction. The fundamental structure of HTVGNN is illustrated in Fig. 1. Firstly, an enhanced temporal perception multi-head self-attention mechanism is proposed to model dynamic temporal correlations. Then, a coupled time-varying graph convolution gated recurrent module (CTVGCRM) is employed to extract spatio-temporal correlations. In the CTVGCRM, we construct static and dynamic time-varying graphs to capture spatial correlations among traffic nodes.
It worth noting that, more accurate temporal multi-attention can provide more differentiated spatio-temporal features for the coupled time-varying graphs. It helps the coupled time-varying graphs to learn traffic patterns at different time steps. In this way, our constructed coupled time-varying graphs can play a better effect. Detailed model structures and mechanisms of our proposed HTVGNN are summarized as follows.
@@ IV-A An enhanced temporal perception multi-head self-attention
where $h$ is the total amount of attention heads. $W$ is the final output projection matrix. However, in traditional multi-head attention mechanism, the key ($K_{i}$) will cause unnecessary interference [17]. In addition, since the learned parameters $W_{i}^{q}$, $W_{i}^{k}$, and $W_{i}^{v}$ remained unchanged after training, the score calculated by the traditional multi-head attention equation defined in Eq. (2) can not fully perceive the temporal trend of self-attention. Therefore, a time-varying mask embedding generator is designed. The key process of the generator finds the daily submask embedding $\tilde{P}^{D}$ and the weekly submask embedding $\tilde{P}^{W}$ of the time corresponding to the current traffic flow $X_{t}$. Then, based on an independent trainable embedding $\tilde{E}\in R^{N\times d_{m}}$, we have proposed an enhanced temporal perception multi-head self-attention. For notational simplicity, our proposed enhanced temporal perception multi-head self-attention (ETPMSA), as shown in Fig.2, can be calculated as follows in Eq.(4):
@@
where $\widehat{W}$ is a learnable parameter, $head_{i}^{ETP}$ denotes the $i^{th}$ self-attention head which can be calculted by the following Eq.(5) to Eq.(9).
@@
where $Q_{i}\in\ R^{T\times N\times d_{q}\ },K_{i}\in\ R^{T\times N\times d_{k}\ }$ and $V_{i}\in\ R^{T\times N\times d_{v}\ }$ are the query, key and value of the traditional time multi-head attention as defined in equation (2), $M_{t}\in R^{N\times N}$ is mask matrix which is defined according to the size of the input data at the time step $t$. It should be noted that, $m_{t}\in R^{N\times d_{m}}$ is the mask embedding which composed of three submask embedding $\tilde{E}\in R^{N\times d_{m}}$ and $\tilde{P}^{D}_{t}\in R^{N\times d_{m}}$ , $\tilde{P}^{W}_{t}\in R^{N\times d_{m}}$. Submask embedding $\tilde{P}^{D}_{t}$ and $\tilde{P}^{W}_{t}$ are the daily submask embedding $\tilde{P}^{D}$ and weekly submask embedding $\tilde{P}^{W}$ at $t$. The submask embedding $\tilde{P}^{D}$ and $\tilde{P}^{W}$ are employed to enhance the temporal perception ability of the model by updating the key $(K)$ of multi-heads self-attention mechanism.
@@ IV-B Position Encodings
Among them, $pos$ represents the position and $d$ represents the feature dimension, $D$ denotes the embedding dimension of the model.
@@ IV-C1 Graph learning
where $E$ is a trainable node embedding, $\ e_{t}$ is a trainable bias of the node embedding $E$. Herein, in order to ensure the diversity of $A_{t}^{s}$ at different time step, we set different bias $e_{t}$ for different time step $t^{th}$ during the training process, $Softmax()$ is used as the activation function.
@@
where $W_{t-T+2},....W_{t}$ are learnable parameter.after completing the graph coupling process,The coupling graph $\tilde{A}_{t}^{s}$ is utilized as the ultimate static adjacency matrix for spatial correlation modeling at each time step.
@@
the dynamic adjacency matrix $\tilde{A}^{v}_{t_{(i,j)}}$ at the $t^{th}$ time step between node $i$ and node $j$ is defined as Eq.(13) to Eq.(16)
@@
Then, Dynamic changes of spatial correlations occuring in the current time should belong to the actual existence and most active nodes. such as nodes which located in adjacency matrix $A$ and DTW matrix $A_{dtw}$. Thus, we combine the adjacency matrix $A$ and DTW matrix $A_{dtw}$ as the mask matrix of dynamic spatial correlations. Finally the dynamic adjacency matrix $\tilde{A}^{v}_{t}$ can be define as Eq.(16):
@@ IV-C2 Coupling time-varying graph convolution gated recurrent unit
where $H_{t}$ and $h_{t}$ are the input hidden embedding and output hidden embedding at the time step $t$, respectively. ||||| | represents the concatenation operation, ${z}_{t}$ and ${r}_{t}$ represent reset gate and update gate the time step $t$, respectively.  $\mathcal{G}_{s}$ and $\mathcal{G}_{d}$ represent the graph convolution network (GCN) and dynamic graph convolution network (DGCN), respectively. $\theta_{z_{1}}$, $\theta_{z_{2}}$ and $\theta_{r_{1}}$, $\theta_{r_{2}}$ and  $\theta_{c_{1}}$, $\theta_{c_{2}}$ are learnable parameters in CTVGCRM.
@@
Dataset PEMS03 PEMS04 PEMS07 PEMS08
@@ V-A Datasets
We validate the predictive performance of our proposed method HTVGNN by conducting experiments on four real-world traffic datasets, namely PEMS03, PEMS04, PEMS07 and PEMS08. These datasets were collected by the California Department of Motor Transportation Performance Measurement System (PeMS). Please refer to Table I for detailed information.
@@ V-B Evaluation Metrics
where $\hat{y}_{i}$ and $y_{i}$ represent the predicted value and real value of traffic flow at node $i$, respectively. $n$ indicates the total number of nodes.
@@ V-D2 Model Settings
The experiments are conducted on a machine equipped with a single NVIDIA RTX3090 and 32GB of memory. HTVGNN is trained using Ubuntu 18.04, PyTorch 2.0.0, and the Adam optimizer [46]. The learning rate is set to 0.001, while the remaining hyperparameters of the model are presented in the table IV.
@@ V-E Performance Comparison
Tables II and III present the comparison results of 12 prediction models on four datasets. Overall, our HTVGNN achieved exceptional prediction accuracy across all 12 prediction steps. The deep learning-based approaches outperformed traditional statistical-based approaches (VAR), indicating their efficacy in modeling highly nonlinear traffic flows. As depicted in Tables II and III, these approaches exhibited suboptimal prediction performance. Moreover, methods that capture spatio-temporal dependencies generally surpassed those solely modeling temporal dependencies (LSTM), underscoring the significance of spatial dependencies. DCRNN and STGCN employ predefined graphs to model spatial dependencies in road networks; however, their performance heavily relies on the quality of these predefined graphs. AGCRN utilizes adaptive graphs to capture spatial dependencies, which typically enhances predictions compared to predefined graphs. Nevertheless, both predefined and adaptive graphs remain static during the prediction stage, disregarding dynamic changes in the road network. In contrast, models that dynamically model spatial dependencies (GMAN, STWave) significantly outperform other approaches based on static graph structures (DCRNN, STGCN, GWN). This can be attributed to their ability to learn dynamically changing graph structures and provide a broader representation space for modeling more complex spatio-temporal dependencies.Compared to other models, HTVGNN incorporates a time-varying mask matrix to enhance the multi-head attention mechanism for accurately capturing dynamic temporal correlations. Moreover, HTVGNN employs a dedicated coupled graph learning method tailored for multi-head attention, effectively improving its performance in long-term prediction tasks. Additionally, to better model dynamic spatial correlations in road networks, we define a mask based on topological and semantic matrices to optimize the dynamic time-varying graphs. Overall, our model outperforms existing state-of-the-art approaches on four datasets, achieving superior prediction accuracy.
@@ V-F Abliation study
$\mathbf{w/o\ TV}$: Replace the time-varying graph in CTVGCRM with a topological graph, emphasizing its pivotal role in capturing spatial correlations.
@@
The experiments were conducted multiple times, and the optimal MAE, RMSE, and MAPE values for each variant on PEMS04 and PEMS08 are presented in Table V.
TABLE V: Ablation experiments on PEMS04 and PEMS08.
@@ V-F2 Effect analysis of different types of graph structure in CTVGCRM and HTVGNN
CTVGCRM: This is a coupling time-varying graph convolution gated recurrent network.
CTVGCRM-sl: Let different time slices be learned with different embeddings, but the spatial correlations learned in different time slices are not coupled.
CTVGCRM-ag: All time slices are learned using the same embedding, in which case the time-varying graph is the same as AGCRN.
CTVGCRM-s: Replace coupled time-varying graphs with topological graphs.
@@
Subsequently, we assess the impact of our proposed enhanced temporal perception multi-head self-attention and coupling time graph on the overall model performance by leveraging the aforementioned optimal mask embedding and dropout techniques. We devise a series of five variants to conduct ablation experiments on our model using PEMS04 and PEMS08 datasets. These variants aim to highlight the significance of time-varying mask embedding and time-varying embedding in our approach, as described below:
@@
$\mathbf{w/o-tm}$: Remove the time-varying mask embedding part of the enhanced temporal perception multi-head self-attention in HTVGNN.
$\mathbf{w/o-cg}$: Remove the coupling part of the time-varying graph in HTVGNN.
$\mathbf{w/o-bc}$: Remove both the time-varying mask embedding part of the enhanced temporal perception multi-head self-attention and the coupling part of the time-varying graph from HTVGNN.
The results presented in Table VIII demonstrate that the superiority of HTVGNN over both w/o-tm, highlighting the positive impact of incorporating time-varying mask embedding on enhancing self-attention accuracy. Furthermore, the superior performance of HTVGNN compared to w/o-cg and the superior performance of w/o-cg compared to w/o-bc indicates its advantage in capturing spatial correlations in transportation networks by leveraging coupling time-varying graphs.
@@
The results presented in Table IX unequivocally demonstrate the superiority of HTVGNN with coupling time-varying graphs over its counterpart without such coupling. This advantage becomes increasingly pronounced as the time step increases. Furthermore, Table IX reveals that HTVGNN outperforms HTVGNN-wc by 4.81%, 4.06%, and 4.29% in terms of MAE, RMSE, and MAPE on PEMS04 dataset at a temporal resolution of 60 minutes, respectively. Similarly, on PEMS08 dataset at the same temporal resolution, HTVGNN achieves improvements of 6.76%, 3.37%, and 6.16% for MAE, RMSE, and MAPE.